Steve and Mimi are placing a border around a rectangular flower bed. One long side will be next to the house and will not have a border. The longer side is to be 3 feet more than the shorter side, and they have 24 feet of border is material. Find the length and width of the flower bed.
Answer by checkley77(12844)

L=S+3
L+2S=24
(S+3)+2S=24
3S=24-3
3S=21
S=21/3
S=7 FT. IS THE SHORT SIDES.
L=7+3=10 FT. FOR THE LONG SIDE.
PROOF:
10+2*7=24
10+14=24
24=24